2x + 3y = 0 x+ 2y = -1
Answer by mananth(16946)

2 x + 3 y = 0 .............1
1 x + 2 y = -1 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 3
-4 x -6 y = 0
3 x + 6 y = -3
Add the two equations
-1 x = -3
/ -1
x = 3
plug value of x in (1)
2 x + 3 y = 0
6 + 3 y = 0
3 y = 0 -6
3 y = -6
y = -2
(3,-2)